# Multiscale scattered data analysis in samplet coordinates

### v1

@@ 4. Samplets for scattered data approximation
uniformly in $\boldsymbol{\alpha},\boldsymbol{\beta}\in\mathbb{N}^{d}$ apart from the diagonal $\Delta\mathrel{\mathrel{\mathop{:}}=}\{({\boldsymbol{x}},{\boldsymbol{y}})\in\Omega\times\Omega:{\boldsymbol{x}}={\boldsymbol{y}}\}$, the matrix ${\boldsymbol{K}}^{\Sigma}$ becomes quasi-sparse and can efficiently be computed. More precisely, there holds the following result, see [7].
@@
where $\tau$ is the cluster supporting $\sigma_{j,k}$ and $\tau^{\prime}$ is the cluster supporting $\sigma_{j^{\prime},k^{\prime}}$, respectively. Then, there exists a constant $C>0$, such that
(19) $\frac{\big{\|}{\boldsymbol{K}}^{\Sigma}-{\boldsymbol{K}}^{\Sigma}_{\varepsilon}\big{\|}_{F}}{\big{\|}{\boldsymbol{K}}^{\Sigma}\big{\|}_{F}}\leq Cm_{q}(r\rho/d)^{-2(q+1)},$
where $m_{q}=\binom{q+d}{d}$ is the dimension of $\mathcal{P}_{q}$ , and the compressed matrix has $\mathcal{O}(m_{q}^{2}N\log N)$ nonzero coefficients.
@@ 5. Multiscale algorithm in samplet coordinates
In this section, we transform the interpolation problem from the original basis, referred to as the natural basis, to the samplet basis, which consists of the scaling distributions and the samplets, as introduced in Section 4. In what follows, we denote the samplet transform for $\mathcal{S}^{\prime}_{\ell}$ by ${\boldsymbol{T}}_{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell}}$ and recall that ${\boldsymbol{T}}^{\intercal}_{\ell}{\boldsymbol{T}}_{\ell}={\boldsymbol{T}}_{\ell}{\boldsymbol{T}}^{\intercal}_{\ell}={\boldsymbol{I}}\in\mathbb{R}^{N_{\ell}\times N_{\ell}}$. Defining the block transform
(20) $\mathcal{T}\mathrel{\mathrel{\mathop{:}}=}\left[\begin{array}[]{cccc}{\boldsymbol{T}}_{1}&&&\\ &{\boldsymbol{T}}_{2}&&\\ &&\ddots&\\ &&&{\boldsymbol{T}}_{L}\end{array}\right],$
there holds $\mathcal{T}^{\intercal}\mathcal{T}=\mathcal{T}\mathcal{T}^{\intercal}={\boldsymbol{I}}$. Hence, the linear system (8) is equivalent to
@@
$\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{cccc}\boldsymbol{K}_{1,1}&&&\\ \boldsymbol{K}_{2,1}&\boldsymbol{K}_{2,2}&&\\ \vdots&\cdots&\ddots&\\ \boldsymbol{K}_{L,1}&\cdots&\boldsymbol{K}_{L,L-1}&\boldsymbol{K}_{L,L}\end{array}\right]\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{c}\boldsymbol{c}_{1}\\ \boldsymbol{c}_{2}\\ \boldsymbol{c}_{3}\\ \vdots\\ \boldsymbol{c}_{L}\end{array}\right]=\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{c}\boldsymbol{f}_{1}\\ \boldsymbol{f}_{2}\\ \boldsymbol{f}_{3}\\ \vdots\\ \boldsymbol{f}_{L}\end{array}\right]$
@@
Letting $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{K}_{\ell,\ell^{\prime}}{\boldsymbol{T}}^{\intercal}_{\ell^{\prime}}$, $\boldsymbol{c}^{\Sigma}_{\ell}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{c}_{\ell}$ and $\boldsymbol{f}^{\Sigma}_{\ell}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{f}_{\ell}$ and dropping the inverse block transform, we arrive at the samplet transformed system
@@
(21) $\left[\begin{array}[]{ccccc}\boldsymbol{K}^{\Sigma}_{1,1}&&&&\\ \boldsymbol{K}^{\Sigma}_{2,1}&\boldsymbol{K}^{\Sigma}_{2,2}&&&\\ \boldsymbol{K}^{\Sigma,3}_{1}&\boldsymbol{K}^{\Sigma}_{3,2}&\boldsymbol{K}^{\Sigma}_{3,3}&&\\ \vdots&\vdots&\cdots&\ddots&\\ \boldsymbol{K}^{\Sigma}_{L,1}&\boldsymbol{K}^{\Sigma}_{L,2}&\cdots&\boldsymbol{K}^{\Sigma}_{L,L-1}&\boldsymbol{K}^{\Sigma}_{L,L}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{c}^{\Sigma}_{1}\\ \boldsymbol{c}^{\Sigma}_{2}\\ \boldsymbol{c}^{\Sigma}_{3}\\ \vdots\\ \boldsymbol{c}^{\Sigma}_{L}\end{array}\right]=\left[\begin{array}[]{c}\boldsymbol{f}^{\Sigma}_{1}\\ \boldsymbol{f}^{\Sigma}_{2}\\ \boldsymbol{f}^{\Sigma}_{3}\\ \vdots\\ \boldsymbol{f}^{\Sigma}_{L}\end{array}\right],$
which is equivalent to (8). In particular, since $\mathcal{T}$ is an isometry, the condition number of the system (21) is identical to that of (8).
@@
Furthermore, we remark that, for a geometric sequence $\{N_{\ell}\}_{\ell}$, the cost for applying $\mathcal{T}$ and $\mathcal{T}^{\intercal}$, respectively, remains linear in $N_{L}$. Each block ${\boldsymbol{K}}^{\Sigma}_{\ell,\ell^{\prime}}$, $\ell\leq\ell^{\prime}$, can be computed with cost $\mathcal{O}(N_{\ell^{\prime}}+N_{\ell}\log N_{\ell})$ if $\{N_{\ell}\}_{\ell}$ forms a geometric sequence, cp. [1, Proof of Theorem 3.16]. Moreover, there are $\mathcal{O}(\log N_{\ell^{\prime}})$ blocks per row, resulting in a computational cost of $\mathcal{O}(N_{\ell^{\prime}}\log N_{\ell^{\prime}})$ per row by the geometric sequence property of $\{N_{\ell}\}_{\ell}$. Summing up the cost for each row, exploiting the geometric sequence again, the overall computation cost for the assembly of the linear system (21) is bounded by $\mathcal{O}(N_{L}\log N_{L})$. To give an idea of the sparsity pattern of the samplet compressed system (21), Figure 1 visualizes the matrix pattern $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}$ using three levels of points $X_{1},X_{2},X_{3}$ arranged on a regular grid for the unit square, with cardinalities of 4 225, 66 049, and 1 050 625, respectively.
@@
Figure 1. Example of the sparsity patterns of the matrices $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}$ for ,$\ell,\ell^{\prime}\in\{1,2,3\}$ in the linear system (21).
@@ 6. Numerical results
We present numerical examples of the proposed multiscale interpolation in samplets coordinates in both two and three dimensional settings. The parameters for the samplet compression are $\rho$, $q$, and $\kappa$. The parameter $\rho$ corresponds to the admissibility condition (18), which steers the a-priorily truncated matrix blocks. A larger $\rho$ results in a larger distance between clusters before they are admissible for compression. In our analysis, we adopt $\rho=d$, so that the bound in Equation (19) is exponentially governed by the number of vanishing moments. This number of vanishing moments corresponds to $q+1$, as specified in (15). The higher $q$, the more polynomial exactness is gained. Lastly, $\kappa$ serves as an a-posteriori compression threshold. After applying the samplet transform to the generalized Vandermonde matrix, we truncate entries smaller than $\kappa$, resulting in a sparser matrix representation.
@@
$\text{error}_{p,{\ell}}=\frac{\|f-s_{\ell}\|_{p}}{\|f\|_{p}}$
for $p=2$ and $p=\infty$ at each of the levels $\ell=1,\ldots,L$. Furthermore, we compute the order of convergence, given by
$\text{order}_{p,{\ell}}=\frac{\log{\left(\displaystyle\frac{\text{error}_{p,{\ell+1}}}{{\text{error}_{p,{\ell}}}}\right)}}{\log{\left(\displaystyle\frac{h_{\ell+1}}{h_{\ell}}\right)}}.$
@@
All computations have been performed at the Centro Svizzero di Calcolo Scientifico (CSCS) on a single node of the Alps cluster [^2] https://www.cscs.ch/computers/alps with two AMD EPYC 7742 @2.25GHz CPUs, up to 420GB of main memory and up to 20 cores. The implementation of the samplet compression is available as software package FMCA [^3] https://github.com/muchip/fmca.

### v2

@@ 5. Samplets for scattered data approximation
uniformly in $\boldsymbol{\alpha},\boldsymbol{\beta}\in\mathbb{N}^{d}$ apart from the diagonal $\Delta\mathrel{\mathrel{\mathop{:}}=}\{({\boldsymbol{x}},{\boldsymbol{y}})\in\Omega\times\Omega:{\boldsymbol{x}}={\boldsymbol{y}}\}$, the matrix ${\boldsymbol{K}}^{\Sigma}$ becomes quasi-sparse and can efficiently be computed. More precisely, a compression based on cluster distances is applied to compute a sparsified version of the matrix ${\boldsymbol{K}}^{\Sigma}$. In detail, there holds the following result, see [9].
@@
where $\tau$ is the cluster supporting $\sigma_{j,k}$ and $\tau^{\prime}$ is the cluster supporting $\sigma_{j^{\prime},k^{\prime}}$, respectively. Then, there exists a constant $C>0$, such that the resulting compressed matrix ${\boldsymbol{K}}^{\Sigma,\rho}$ satisfies
(19) $\frac{\big{\|}{\boldsymbol{K}}^{\Sigma}-{\boldsymbol{K}}^{\Sigma,\rho}\big{\|}_{F}}{\big{\|}{\boldsymbol{K}}^{\Sigma}\big{\|}_{F}}\leq Cm_{q}\bigg{(}\frac{r\rho}{d}\bigg{)}^{-2(q+1)},$
where $m_{q}=\binom{q+d}{d}$ is the dimension of $\mathcal{P}_{q}$. The compressed matrix has $\mathcal{O}(m_{q}^{2}N\log N)$ nonzero coefficients.
We remark that, given a radial function satisfying (17), the corresponding scaled kernel (6) satisfies the asymptotical smoothness estimate, cp. [10],
$\bigg{|}\frac{\partial^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}}{\partial{\boldsymbol{x}}^{\boldsymbol{\alpha}}\partial{\boldsymbol{y}}^{\boldsymbol{\beta}}}K_{\ell}({\boldsymbol{x}},{\boldsymbol{y}})\bigg{|}\leq C\delta_{\ell}^{-d}\frac{(|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|)!}{(r\delta_{\ell})^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}\|({\boldsymbol{x}}-{\boldsymbol{y}})/\delta_{\ell}\|_{2}^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}}=C\delta_{\ell}^{-d}\frac{(|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|)!}{r^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}\|{\boldsymbol{x}}-{\boldsymbol{y}}\|_{2}^{|\boldsymbol{\alpha}|+|\boldsymbol{\beta}|}},$
which means that the dependency of the derivatives on $\delta_{\ell}$ just cancels out. Moreover, we remark that, since we only require derivatives of the kernel up to order $12q+$ 1 to be bounded, the asymptotical smoothness requirement could in principle also be weakened to kernels of finite smoothness, see, e.g., [11] for details.
@@ 6. Multiscale algorithm in samplet coordinates
In this section, we transform the interpolation problem from the original basis, referred to as the kernel basis, to the samplet basis, as introduced in Section 5. In what follows, we denote the samplet transform for $\mathcal{S}^{\prime}_{\ell}$ by ${\boldsymbol{T}}_{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell}}$ and recall that ${\boldsymbol{T}}^{\intercal}_{\ell}{\boldsymbol{T}}_{\ell}={\boldsymbol{T}}_{\ell}{\boldsymbol{T}}^{\intercal}_{\ell}={\boldsymbol{I}}\in\mathbb{R}^{N_{\ell}\times N_{\ell}}$. Defining the block transform $\mathcal{T}\mathrel{\mathrel{\mathop{:}}=}\operatorname{diag}({\boldsymbol{T}}_{1},\ldots,{\boldsymbol{T}}_{L})$ there holds $\mathcal{T}^{\intercal}\mathcal{T}=\mathcal{T}\mathcal{T}^{\intercal}={\boldsymbol{I}}$. Hence, the linear system (10) is equivalent to
@@
$\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{cccc}\boldsymbol{K}_{1,1}&&&\\ \boldsymbol{K}_{2,1}&\boldsymbol{K}_{2,2}&&\\ \vdots&\vdots&\ddots&\\ \boldsymbol{K}_{L,1}&\boldsymbol{K}_{L,2}&\cdots&\boldsymbol{K}_{L,L}\end{array}\right]\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{c}\boldsymbol{c}_{1}\\ \boldsymbol{c}_{2}\\ \vdots\\ \boldsymbol{c}_{L}\end{array}\right]=\mathcal{T}^{\intercal}\mathcal{T}\left[\begin{array}[]{c}\boldsymbol{f}_{1}\\ \boldsymbol{f}_{2}\\ \vdots\\ \boldsymbol{f}_{L}\end{array}\right].$
@@
Letting $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{K}_{\ell,\ell^{\prime}}{\boldsymbol{T}}^{\intercal}_{\ell^{\prime}}\in\mathbb{R}^{N_{\ell}\times N_{\ell}^{\prime}},\ \boldsymbol{c}^{\Sigma}_{\ell}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{c}_{\ell}\in\mathbb{R}^{N_{\ell}},\ \boldsymbol{f}^{\Sigma}_{\ell}\mathrel{\mathrel{\mathop{:}}=}{\boldsymbol{T}}_{\ell}\boldsymbol{f}_{\ell}\in\mathbb{R}^{N_{\ell}},$ we can rewrite the previous system according to
@@
(20) $\left[\begin{array}[]{cccc}\boldsymbol{K}^{\Sigma}_{1,1}&&&\\ \boldsymbol{K}^{\Sigma}_{2,1}&\boldsymbol{K}^{\Sigma}_{2,2}&&\\ \vdots&\vdots&\ddots&\\ \boldsymbol{K}^{\Sigma}_{L,1}&\boldsymbol{K}^{\Sigma}_{L,2}&\cdots&\boldsymbol{K}^{\Sigma}_{L,L}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{c}^{\Sigma}_{1}\\ \boldsymbol{c}^{\Sigma}_{2}\\ \vdots\\ \boldsymbol{c}^{\Sigma}_{L}\end{array}\right]=\left[\begin{array}[]{c}\boldsymbol{f}^{\Sigma}_{1}\\ \boldsymbol{f}^{\Sigma}_{2}\\ \vdots\\ \boldsymbol{f}^{\Sigma}_{L}\end{array}\right].$
Since $\mathcal{T}$ is an isometry, the condition number of this system is identical to that of (10). In particular, the diagonal blocks remain well-conditioned in accordance with Corollary 3.2. Applying the matrix compression from Theorem 5.2 to each of the blocks $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}$, we arrive at the compressed, samplet transformed system
(21) $\left[\begin{array}[]{cccc}\boldsymbol{K}^{\Sigma,\rho}_{1,1}&&&\\ \boldsymbol{K}^{\Sigma,\rho}_{2,1}&\boldsymbol{K}^{\Sigma,\rho}_{2,2}&&\\ \vdots&\vdots&\ddots&\\ \boldsymbol{K}^{\Sigma,\rho}_{L,1}&\boldsymbol{K}^{\Sigma,\rho}_{L,2}&\cdots&\boldsymbol{K}^{\Sigma,\rho}_{L,L}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{c}^{\Sigma}_{1}\\ \boldsymbol{c}^{\Sigma}_{2}\\ \vdots\\ \boldsymbol{c}^{\Sigma}_{L}\end{array}\right]=\left[\begin{array}[]{c}\boldsymbol{f}^{\Sigma}_{1}\\ \boldsymbol{f}^{\Sigma}_{2}\\ \vdots\\ \boldsymbol{f}^{\Sigma}_{L}\end{array}\right].$
A detailed algorithm for solving (21) using block forward substitution and evaluating the resulting approximation at a set of points can be found in Algorithm (1). We remark that replacing the kernel matrix for the evaluation points by its samplet compressed version leads to a further consistency error, which can, however, be bounded with similar ideas as in Lemma 4.1.
Algorithm 1 Multiscale interpolation in samplet coordinates
Input: Sets of points $\{X_{1},\ldots,X_{L}\}$, function $f$, parameters $\rho,q$.
Output: Coefficient vectors $\widehat{\boldsymbol{c}}_{1},\ldots,\widehat{\boldsymbol{c}}_{L}$ corresponding to $\widehat{w}_{1},\ldots,\widehat{w}_{L}$.
1:  for $\ell=1,2,\ldots,L$ do
2:     Compute ${\boldsymbol{f}}^{\Sigma}_{\ell}\leftarrow{\boldsymbol{T}}_{\ell}{\boldsymbol{f}}_{\ell}\mathrel{\mathrel{\mathop{:}}=}f|X_{\ell}$.
3:     for $\ell^{\prime}=1,2,\ldots,\ell$ do
4:        Compute $\boldsymbol{K}^{\Sigma,\rho}_{\ell,\ell^{\prime}}$ using $q+1$ vanishing moments.
5:     end for
6:     Compute $\widehat{\boldsymbol{e}}^{\Sigma}_{\ell}\leftarrow\boldsymbol{f}^{\Sigma}_{\ell}-\sum_{\ell^{\prime}=1}^{\ell-1}\boldsymbol{K}^{\Sigma,\rho}_{\ell,\ell^{\prime}}\widehat{\boldsymbol{c}}^{\Sigma}_{\ell^{\prime}}$.
7:     Solve the local interpolation problem $\boldsymbol{K}^{\Sigma,\rho}_{\ell,\ell}\widehat{\boldsymbol{c}}^{\Sigma}_{\ell}=\widehat{\boldsymbol{e}}^{\Sigma}_{\ell}$.
8:     Compute $\widehat{\boldsymbol{c}}_{\ell}\leftarrow{\boldsymbol{T}}_{\ell}^{\intercal}\widehat{\boldsymbol{c}}^{\Sigma}_{\ell}$.
9:  end for
The overall compression error of (21), can be estimated by summing up the block-wise errors in accordance with Theorem 5.2. The following result is immediate.
Theorem 6.1.
Let $X_{1},\ldots,X_{L}$ be quasi-uniform. Then, the compression error in the linear system (21) can be bounded according to
$\frac{\big{\|}{\mathcal{K}}^{\Sigma}-{\mathcal{K}}^{\Sigma,\rho}\big{\|}_{F}}{\big{\|}{\mathcal{K}}^{\Sigma}\big{\|}_{F}}\leq Cm_{q}(r\rho/d)^{-2(q+1)},$
where we denote the block matrix from (20) by ${\mathcal{K}}^{\Sigma}$ and the one from (21) by ${\mathcal{K}}^{\Sigma,\rho}$, respectively.
Observing that $\mathcal{T}\mathcal{D}^{1/2}=\mathcal{D}^{1/2}\mathcal{T}$, combining the previous theorem with Theorem 4.6 immediately results in an error estimate for the approximation $\widehat{s}_{L}$ to $s_{L}$ obtained by solving the samplet compressed system.
Theorem 6.2.
Let the conditions of Theorem 4.6 be satisfied. Then, the numerical approximation $\widehat{s}_{L}$ resulting from solving (21) satisfies the error estimate
$\|f-\widehat{s}_{L}\|_{L^{2}(\Omega)}\leq C\Big{(}(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{L}+\sqrt{L}m_{q}(r\rho/d)^{-2(q+1)}\Big{)}\|f\|_{H^{\theta}(\Omega)}.$
Finally, we have the subsequent result on the cost of our multiscale algorithm, which is in line with a similar result for multiscale interpolation using hierarchical matrices, see [14].
Theorem 6.3.
Let $X_{1},\ldots,X_{L}$ be quasi-uniform and assume that the corresponding cardinalities $\{N_{\ell}\}_{\ell}$ form a geometric sequence, i.e., $N_{1}>0$ and $N_{\ell}=c_{\mathrm{geo}}^{\ell}N_{1}$ for some constant $c_{\mathrm{geo}}>1$. Then, the overall computational cost for assembling the linear system (21) and the number of non-vanishing matrix entries are both bounded by $\mathcal{O}(N_{L}\log^{2}N_{L})$, with the hidden constant depending on the dimension $m_{q}$ of $\mathcal{P}_{q}$.
Proof.
Fix a row index $\ell$. It has been shown within [1, Proof of Theorem 3.16] that the cost for the computation of $\mathbf{K}^{\Sigma,\rho}_{\ell,\ell^{\prime}}$ amounts to $\mathcal{O}\left(N_{\ell}\log N_{\ell^{\prime}}\right)$, which is also the number of non-vanishing entries of that row. Moreover, there are $\ell$ blocks per row. Hence, the cost for the computation of the $\ell$-th row of (21) is given by
$\sum_{\ell^{\prime}=1}^{\ell}\mathcal{O}\left(N_{\ell}\log N_{\ell^{\prime}}\right)=\mathcal{O}\bigg{(}N_{\ell}\sum_{\ell^{\prime}=1}^{\ell}\log N_{\ell^{\prime}}\bigg{)}=\mathcal{O}\left(N_{\ell}\log^{2}N_{\ell}\right),$
since $\ell=\mathcal{O}(\log N_{\ell})$ due to the geometric sequence property. Next, summing up the cost of the individual rows yields the overall cost
$\sum_{\ell=1}^{L}\mathcal{O}\left(N_{\ell}\log^{2}N_{\ell}\right)=\mathcal{O}\bigg{(}\sum_{\ell=1}^{L}N_{\ell}\log^{2}N_{\ell}\bigg{)}=\mathcal{O}\bigg{(}\log^{2}N_{L}\sum_{\ell=1}^{L}N_{\ell}\bigg{)}=\mathcal{O}\left(N_{L}\log^{2}N_{L}\right),$
again due to the geometric sequence property, i.e., $\sum_{\ell=1}^{L}N_{L}=\mathcal{O}(N_{L})$. This completes the proof. ∎
@@
To give an idea of the sparsity pattern of the samplet compressed system (21), Figure 1 visualizes the matrix pattern $\boldsymbol{K}^{\Sigma}_{\ell,\ell^{\prime}}$ using three levels of points $X_{1},X_{2},X_{3}$ arranged on a regular grid for the unit square, with cardinalities of 16 641, 66 049, and 263 169, respectively. Furthermore, we remark that, for a geometric sequence $\{N_{\ell}\}_{\ell}$, the cost for applying $\mathcal{T}$ and $\mathcal{T}^{\intercal}$, respectively, remains linear in $N_{L}$.
@@
Figure 1. Example of the sparsity patterns of the matrices $\boldsymbol{K}^{\Sigma,\rho}_{\ell,\ell^{\prime}}$ for ,$\ell,\ell^{\prime}\in\{1,2,3\}$ of the compressed linear system from (21).
@@ 7. Numerical results
We present numerical examples of the proposed multiscale interpolation in samplets coordinates in both two and three dimensional settings. The parameters for the samplet compression are $\rho$ and $q$. The first one is associated with the admissibility condition (18), which determines the matrix blocks that are initially truncated. A larger $\rho$ results in a larger distance between clusters before they are admissible for compression. In our analysis, we adopt $\rho=d$, so that the bound in Equation (19) is exponentially governed by the number of vanishing moments, denoted as $q+1$, see (15). The higher $q+1$, the more polynomial exactness is gained.
In the experiments, we vary also the parameter $\gamma$, defined as $\gamma=\eta\mu,$ where $\delta$ is the kernel rescaling factor, given by $\delta_{i}=\eta h_{X_{i}}$ and previously defined in (3), while $\eta$ represents the scaling factor for the fill distance between consecutive levels, i.e., $h_{X_{\ell+1}}\approx\mu h_{X_{\ell}}$.
@@
$\text{error}_{p,{m}}=\frac{\|f-\widehat{s}_{m}|X_{\text{eval}}\|_{\ell^{p}}}{\|f|X_{\text{eval}}\|_{\ell^{p}}}$
for $p=2$ and $p=\infty$ at each of the levels $m=1,\ldots,L$ for a given set of evaluation points $X_{\text{eval}}$. Furthermore, we compute the order of convergence, given by
$\text{order}_{p,{m}}=\frac{\log{\left(\displaystyle\frac{\text{error}_{p,{m+1}}}{{\text{error}_{p,{m}}}}\right)}}{\log{\left(\displaystyle\frac{h_{m+1}}{h_{m}}\right)}}.$
@@
All computations have been performed at the Centro Svizzero di Calcolo Scientifico (CSCS) on a single node of the Alps cluster [^2] https://www.cscs.ch/computers/alps with two AMD EPYC 7742 @2.25GHz CPUs with 470GB of main memory. For the computations, we have used up to 16 cores. The implementation of the samplet compression is available as software package FMCA [^3] https://github.com/muchip/fmca.